OpenStudy (anonymous):

medals an fan!!!!!!!!!!!!!!!!!!!!!!!!!Which number would you add to three and nine-tenths to give an approximate sum of 6? two and one-third three and two-elevenths one and one-twelfth two and seven-eighths

jimthompson5910 (jim_thompson5910):

let x be that unknown number

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

x + 3.9 = 6 to solve this for x, we subtract 3.9 from both sides x + 3.9 - 3.9 = 6 - 3.9 x + 0 = 6 - 3.9 x + 0 = 2.1 x = 2.1

jimthompson5910 (jim_thompson5910):

So that unknown number you add on is 2.1

OpenStudy (anonymous):

huh @jim_thompson5910

jimthompson5910 (jim_thompson5910):

where are you stuck?